Find g(f(x)) for the following:
f(x) = (x^3) - 1, g(x) = (x+1)^(1/3)

Answer by venugopalramana(3286)

Find g(f(x)) for the following:
LET Y=f(x) = (x^3) - 1,
THEN G(F(X))=G(Y)
g(x) = (x+1)^(1/3)
G(Y)=(Y+1)^(1/3) = [X^3-1+1]^(1/3)]=[X^3]^(1/3)=X^(3*1/3)=X